OpenStudy (anonymous):

Write an equation in point-slope form of the line that passes through (5, 0), and whose slope is -1/2. Help Please!!!

OpenStudy (anonymous):

K so the point-slope formula is y-y1=m(x-x1). M= Slope, and (x1,y1) for the coords. When we put this in we get y-0=-1/2(x-5) And it's in point-slope formula right now :)

OpenStudy (anonymous):

Thank you so much!